ncome statements and balance sheets follow for Microsoft Corporation and Apple Inc. Refer to these financial statements to answer the requirements. MICROSOFT CORPORATION Income Statements For the years ended June 30, (in millions) Revenue 2016 2015 $75,956 17 624 Product $61,502 23,818 85,320 Service 93,580 Total revenue Cost of revenue 17,880 14,900 32,780 52,540 11,988 14,697 4,563 1,110 20,182 903 (1,243) 91 19,751 2,953 $16,798 21,410 11,628 33,038 60,542 12,046 15,713 4,611 10,011 18,161 766 Product Service and other Total cost of revenue Gross margin Research and development Sales and marketing General and administrative Impairment, integration, and restructuring Operating income Dividends and interest income Interest expense Other income (expense), net Income before taxes Provision for income taxes Net income (781) 361 18,507 6,314 $ 12,193 Continued next page
media%2F3d8%2F3d8dee86-49d1-4ee7-84c7-1e
media%2Ff3b%2Ff3b0e773-4d85-4288-be05-be
media%2F0a4%2F0a49d2da-3495-4b7e-b233-76
Answer #1

Solition:

calculations for both companies at 2016.

Ratio Formula Calculation MICROSOFT Calculation APPLE
Current ratio: Current asset / Current liabilities 139,660 / 59,357 2.35 106,869 / 79,006 1.35
Quick ratio: Current assets - inventories - prepaid expenses / Current liabilities - Bank overdraft 139,660 - 2251 / 59,357 2.31 106,869 / 79,006 1.32
Times interest earned ratio : Devidends and Interest income / net income 903 / 16,798 0.05 3,999 / 45,687 0.09
Liquidity to equity ratio: Current liabilities + loan + long term liabilities / Share capital + Retained earnings 59,357 + 40,783 + 13,640 / 68,178 + 2,282 + 1,537 = 113,780 / 71997 1.58 79,006 + 75,427 + 36,074 / 31,251 + 96,364 + 634 = 190,507 / 128,249 1.48

Explaination:

(a). Microsoft corporation is more liquid on its current ratio and quick ratio is higher than apple inc.

(b). Apple inc debt equity ratio is less than microsoft which means it in less dependable on outside fund. So, it is more solvent.

(c). Both companies have debt equity ratio is greater than one. Which means they have investment more from outside parties.